⋆⋆  ⋆⋆ Pégase - HR

Projet d’Étude des GAlaxies par Synthèse Évolutive - Haute Résolution

Version 1.1: January 19, 2005

Damien Le Borgne — Michel Fioc — Brigitte Rocca-Volmerange Philippe Prugniel — Ariane Lançon — Caroline Soubiran

Institut d’astrophysique de Paris 98 bis boulevard Arago, 75014 Paris, France

E-mail: leborgne@iap.fr, fioc@iap.fr
WWW: http://www.iap.fr/pegase/

Contents

1 Introduction 3

2 Installation 3 2.1 Structure of the package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Basic installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 Test of the installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.4 Organization of the files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.5 System requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 Differences between PÉGASE.2 and PÉGASE-HR 6 3.1 Stellar libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 Algorithms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.3 Changes in SSPs files and scenarios files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.4 FITS outputs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Usage of PÉGASE-HR 8 4.1 Computing isochrones: SSPs HR.f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.2 Describing scenarios of evolution: scenarios HR.f . . . . . . . . . . . . . . . . . . . . . . . 8 4.3 Computing synthetic spectra: spectra HR.f . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.4 Measuring Lick indices on high-resolution synthetic spectra . . . . . . . . . . . . . . . . . 9 4.5 Measuring colors on low-resolution synthetic spectra . . . . . . . . . . . . . . . . . . . . . 9

5 References 10

1 Introduction

PÉGASE-HR (presented in Le Borgne et al. 2004) is a code aimed at computing synthetic evolutionary optical spectra of galaxies with a very high resolution (R=10 000, or δλ = 0.55 Å) over the wavelength range λ = [4000, 6800] Å. It’s the result of combining the code PÉGASE.2 with the high-resolution stellar library ÉLODIE. This code can also be used at low resolution (R ≃ 200) over the range covered by the BaSeL library (from far UV to the near IR), and then produces the same results as PÉGASE.2.

The PÉGASE.2 code (Version 2 of PÉGASE, Projet d’Etude des GAlaxies par Synthese Evolutive in French) is aimed at modeling the spectral evolution of galaxies. PÉGASE.2 is presented in Fioc & Rocca-Volmerange (1997) and the code is fully described in a README file (Fioc & Rocca-Volmerange

2000) (ftp://ftp.iap.fr/pub/from users/pegase/PEGASE.2/, also available in the doc/ directory of

this package). PÉGASE.2 uses the BaSeL (Lejeune et al. 1997,1998) library of stellar spectra and can therefore synthesize low-resolution (with a resolution R ≃ 200, from 91 Å to 160 μm) ultraviolet to near-infrared spectra of Hubble sequence galaxies as well as of starbursts.

In PÉGASE-HR, the BaSeL library can be replaced by a grid of spectra interpolated from the highresolution ÉLODIE library of stellar spectra.

The ÉLODIE library is a stellar database of 1959 spectra for 1503 stars, observed with the echelle spectrograph ÉLODIE on the 193 cm telescope at the Observatoire de Haute Provence. Previous versions of the library are presented in Prugniel & Soubiran (2001) and Soubiran et al. (1998); Katz et al. (1998). It has been updated for the present work by doubling the number of spectra, which greatly improved the coverage of the parameter space (in effective temperature, surface gravity, and metallicity). The data reduction has also been improved, in particular the flux calibration, and the wavelength range has been extended to 400-680 nm. For the purpose of population synthesis, the original resolution R=42 000 has been reduced to R=10 000 at lambda=550 nm, or more precisely to a gaussian instrumental profile of FWHM≃ 0.55 Å over the whole range of wavelengths. The typical signal-to-noise (S/N) ratio of the spectra is 500/Å. The HR diagram coverage with the estimated stellar parameters is extensive (see Le Borgne et al. 2004 for details). The up-to-date version of the library of stellar spectra is available at

http://www.obs.u-bordeaux1.fr/public/astro/CSO/elodie library.html where fully reduced spec-

tra, as well as the estimated stellar parameters, are provided. From this library, a grid of stellar spectra with regularly spaced stellar parameters (the grid has the same underlying shape as the Lejeune et al. grid) was created by interpolation. This grid (stellibELODIE.fits), available at http://www.iap.fr/pegase/, is used by PÉGASE-HR to compute synthetic spectra. This document describes the installation procedure, the main changes between PEGASE.2 and PEGASE-HR’s algorithms and output formats, and a simplified description of the steps one should follow to compute high resolution synthetic spectra.

2 Installation

2.1 Structure of the package

Before the installation, the package contains the following files:

  ./configure
  ./src/                        ! Source code
  ./doc/                        ! Documentation
  ./data/                       ! Data files
     ./data/tracks/                 -> stellar tracks
     ./data/stellibs/               -> stellar libraries
                                       (stellibLCBcor.fits and stellibELODIE.fits)
     ./data/external/               -> various data files
     ./data/tests/                  -> files used to test the installation
     ./data/user_defined/           -> data files that can be modified by
                                       the user: ages.dat, IMFs, etc.
     ./data/isochrones/             -> pre-computed isochrones (with SSPs_HR.f)

If the ÉLODIE library data/stellibs/stellibELODIE.fits is not in the package, you should download it from http://www.iap.fr/pegase/ and copy it into the directory data/stellibs/. Warning: the size of this file is 150 Mo. After the installation, the additionnal directories should be present:

  ./bin/                        ! Binary files
  ./lib/                        ! Binary libraries

2.2 Basic installation

The compilation and installation is made by the standard linux commands:

      > ./configure
      > make

Then:

      > source peg_cshrc       (if your shell is csh or tcsh)
    OR
      > . peg_bashrc           (if your shell is bash)

Details of the installation:

• configure : ./configure generates the file Makefile. The presence of the required cfitsio library is checked during this process. If this library is not installed on your computer, you can download

     it from http://heasarc.gsfc.nasa.gov/docs/software/fitsio/fitsio.html and install it. If

the library is installed but not found, please use the argument “--with-libraries=” in your call to ./configure to specify the path where the library can be found.

     Example: > ./configure --with-libraries=/usr/lib/

• make : make uses the information in Makefile to compile the various parts of the PÉGASE-HR code. The following binary files are built:

       – bin/SSPs HR
       – bin/scenarios HR

       – bin/spectra HR
       – bin/lick
       – bin/fitstodat

• Setting up the PATH variable: Depending on the shell you use, you can use the scripts peg bashrc and peg cshrc which are created in the “make” process. These scripts add the bin/ directory to the environment variable PATH.

2.3 Test of the installation

Once the package is installed, and once you have downloaded the ÉLODIE grid stellibELODIE.fits and copied it into the data/stellib/ directory, you can test the results of your spectra HR compiled code by running

   > make test

This will compute three scenarios and measure Lick indices on a high resolution spectrum, in the temporary directory test tmp. Then it will print the result of the comparison of pre-computed spectra (stored in data/tests/) with the spectra computed on your machine.

2.4 Organization of the files

Unlike PÉGASE.2, it is not required to run the PÉGASE-HR codes from the directory containing the data files. During the installation, the codes are compiled in such a way that the data files are always read in the data/ directory. The data/ directory is divided into several sub-directories (see the data/README file for more details). The commonly used data files that the user can modify are in the data/user defined directory. The most important files that can be modified (but should not be deleted) are:

ages.dat            ! list of the output ages for the spectra (in Myr)
filters.dat         ! definition of the filters' transmission curves
index_table.dat     ! definition of the Lick indices
list_IMFs.dat       ! list of the IMFs to choose from (when using SSPs_HR)
list_tracks.dat     ! list of the stellar tracks to use (when calling SSPs_HR)
list_stellibs.dat   ! list of the stellar libraries to use
                                             (used by SSPs_HR and spectra_HR)

2.5 System requirements

• Memory : To run the code spectra HR, you need at least 200 Mo of memory. If necessary, you might want to loosen the system limits by running the following commands:

          > unlimit
          > limit coredumpsize 0

• Disk space : you will need a minimum of 70 Mo of space disk for package. This does not include the ÉLODIE stellar library (stellibELODIE.fits) which, alone, is 150 Mo large.

3 Differences between PÉGASE.2 and PÉGASE-HR

This section will be useful to users already familiar with PÉGASE.2, and may be skipped by others readers.

3.1 Stellar libraries

PÉGASE-HR is designed to run with a variety of stellar spectral libraries, and two libraries are provided within the distributed package. The use of a FITS format for these libraries speeds up access to the data considerably compared to the previous ASCII format of PÉGASE.2. The header keyword grid typ defines the grid for which stellar spectra are available (a grid point is characterized by its effective temperature, surface gravity and metallicity). grid typ takes the value standard for the library of Lejeune et al. 1998, in the format adopted for PÉGASE.2 (stellibLCBcor.fits). It also takes the value standard for any other library provided on the same grid. The ÉLODIE grid (stellibELODIE.fits) contains fewer spectra than the BaSeL one (Lejeune et al. 1998). For this grid, grid typ takes the value truncated. The output of SSPs HR will be identical for all stellar libraries that have the same parameter grid (and thus the same grid typ).

3.2 Algorithms

Although the program sources have been reorganized, most algorithms of PÉGASE-HR are identical to those of PÉGASE.2. The sequence of programs to run in order to produce a database of single stellar populations, define a star formation scenario, and compute the corresponding galaxy spectrum, is also unchanged.

PÉGASE.2 PÉGASE-HR
SSPs.f  SSPs HR.f
scenarios.f scenarios HR.f
spectra.f  spectra HR.f

When used with the low resolution stellar library BaSeL (stellibLCBcor), PÉGASE-HR produces the same synthetic spectra as PÉGASE.2, though in a different format (cf. Sect. 3.4). The ÉLODIE library being less extended (in metallicity, effective temperature and surface gravity) than the BaSeL library, the algorithm of SSP HR.f had to be modified from PÉGASE.2’s version. In SSPs.f, the interpolation of the stellar spectra made use of the spectra with metallicities immediately greater and smaller than the metallicity of the tracks. In SSPs HR.f, when no spectra with metallicities close to the metallicity of the tracks are available in the ÉLODIE stellar library, some spectra with metallicities larger or smaller are used in the interpolation. A new code (lick.f) can be used to measure some Lick indices on the evolutionary spectra (see Sect. 4.4 for details.)

3.3 Changes in SSPs files and scenarios files

The use of stellar libraries with different grids makes it necessary to secure the use of SSPs files and of stellar libraries. This is done by the keyword grid typ, as described above. The files prefix SSPs.dat now contain a header specifying the type and the name of the grid used to compute the isochrones. This header is produced automatically by the code SSPs HR. The individual isochrone files (e.g. prefix tracksZ0.02.dat) now also contain in their header an additionnal line specifying the type of the grid used. The code scenarios HR now adds a third line in the header of the scenario files, with the name of the stellar library that will be used to compute the synthetic spectra. The code spectra HR then checks the compatibility between the grid used to compute the isochrones (the files prefix SSP tracks*.dat) and the grid defined in the scenario, used to compute the synthetic spectrum.

3.4 FITS outputs

To reduce the size of the output files, and to increase the speed of the access to the models, the outputs of the code spectra HR are now in binary FITS format. The structure of the output FITS files is described here. The FITS files are made of several blocks (or extensions, or HDU for Header Data Units).

• the primary extension contains, in a binary table, the evolutionary spectra. The first dimension is the number of wavelengths; the second dimension is the number of timesteps (hereafter ntimes). The spectra are still normalized to one solar mass of galaxy, but they are now in units of L_{⊙}.Å^{−1} (with L_{⊙} = 3.826 × 10^{33} erg.s^{−1}), instead of erg.s^{−1}.Å^{−1} in PÉGASE.2.

• the extension ETS LINES contains the emission lines computed for a typical Hii region, following the prescriptions used in PÉGASE.2. It has 2 columns and 61 rows (61 being the number of emission lines computed). The first column contains the wavelengths of the lines. The second column contains, at each row, ntimes values, i.e. the evolution of the flux for each emission line.

• the extension ETS PARA contains the evolution of all the parameters for the galaxy (masses, metallicities, extinction, rates of supernovae Ia and II, number of Lyman photons, etc.) as described in PÉGASE.2’s README file.

• the extension ETS CONT WCA, if present, contains a vector of the wavelengths for which the synthetic spectra are given. This extension is omitted if the wavelength scale is linearly regular (which is the case for the ÉLODIE library). In this case, the wavelength scale is fully described in the header of the primary extension by the reference wavelength (keyword CRVAL1) and the wavelength step (keyword CDELT1).

• the extension ETS LICK, if present, contains the values of the Lick indices (as defined in the file data/user defined/index table.dat) measured on the synthetic evolutionary spectra. The first dimension of the binary table is the number of indices; the second dimension is the number of timesteps. This extension is added to the FITS file by the code lick.f.

The code fitstodat included in the package makes it possible to convert a PÉGASE-HR FITS output file to an ASCII file, in the same format as PÉGASE.2’s outputs. This code can be useful for people used to PÉGASE.2’s outputs and who do not want to change the reading format of their own routines. However, we stress that the ASCII files for PÉGASE-HR outputs are about ten times as large as PÉGASE.2’s output files (i.e. approximately 10 Mo, depending on the number of timesteps). Note also that the information on Lick indices, if present, is lost during the conversion.

4 Usage of PÉGASE-HR

Computing spectra with PÉGASE-HR is very similar to the way to do it with PÉGASE.2. If you are not familiar with PÉGASE.2, we advise you strongly to read the PÉGASE.2’s README file (included in the doc/ directory). We describe here the simple steps one should follow to compute a synthetic spectrum with PÉGASE-HR.

4.1 Computing isochrones: SSPs HR.f

The first step is to compute isochrones for various metallicities. Run

    > SSPs_HR

and make your choices of IMF, model for SNe ejecta, stellar winds, and stellar library. PÉGASE.2’s README file describes the way to add other IMFs. The stellar libraries available by default are:

• stellibLCB.fits : Lejeune et al. (1998) low-resolution stellar library

• stellibELODIE.fits : ÉLODIE (high-resolution) stellar library interpolated onto a regular grid of stellar parameters.

A file (prefix SSPs.dat) is then created, and contains the list of the isochrone files for every metallicities.

The isochrone files (prefix tracksZ0.0001.dat . . . prefix tracksZ0.1.dat) are also created in the

current directory. The metallicities of the isochrones computed with the ÉLODIE library can, in principle, be the same as the metallicities of the PÉGASE.2 isochrones. However, the small number of stars with extreme metallicities in this library makes the isochrones with very low or very large metallicities more uncertain.

4.2 Describing scenarios of evolution: scenarios HR.f

You can now define one or several scenarii of evolution, from the simplest (single instantaneous burst) to the most complex (with infall, galactic winds, continuous star formation, evolution of the metallicity, etc.). Run

    > scenarios_HR

and choose

• the name of the text file created by this code (the scenario itself)

• the name of the file containing the list of isochrones (prefix SSPs.dat). A default name is given, and corresponds to a set of isochrones computed with a Salpeter IMF, stellar winds, and the ÉLODIE library. In the next step (spectra HR), the isochrones will be read automatically in the same directory as prefix SSPs.dat.

• the fraction of sub-stellar objects

• the stellar library you want to use. Warning: the stellar library should be the same as the one used to create prefix SSPs.dat. If it is not the case, an error will appear in the next step (spectra HR) and no synthetic spectrum will be computed.

• for each scenario of evolution, the parameters of star formation, as described in details in PÉGASE.2’s README.

4.3 Computing synthetic spectra: spectra HR.f

Once a scenario is defined (and stored in a text file, for example myscenario.scn), run spectra HR to compute the evolutionary synthetic spectra:

    > spectra_HR myscenario.scn

If you omit the command-line argument (here, myscenario.scn), you will be asked for the name of the file describing the scenarios. One or several output files (in FITS format, with the structure described in section 3.4) are then created. If these files already exist, the program will abort, issuing an error message. Important notice: if a file ages.dat is present in the current directory, then the output spectra are computed for the timesteps listed in this particular file, overriding the default ages.dat (located in the PÉGASE-HR directory data/user defined/). In this case, a warning is given during the execution of

spectra HR.

4.4 Measuring Lick indices on high-resolution synthetic spectra

Once high-resolution evolutionary spectra are computed, one can measure the classic Lick indices by running for example

    > lick myspectrum.fits 5.

The first argument is the evolutionary synthetic spectrum; the second argument is the FWHM of the emission lines. In the example above, the predicted emission lines (contained in the extension ETS LINES of the FITS file) are added to the spectrum, with FWHM=5 Å for each line, before measuring the indices. The emission lines are computed as in PÉGASE.2 for a mean Hii region, consistently with the ionizing flux of the massive stars. If one doesn’t want to include the emission lines in the measurement, they should set FWHM=0 (by

running lick spectrum 0.).

If lick is called without any command-line argument, then the user is prompted for the filename and the FWHM.

4.5 Measuring colors on low-resolution synthetic spectra

When low resolution synthetic spectra are computed with PÉGASE-HR, one can measure the colors on these spectra, just like in PÉGASE.2. The code to measure the colors is colors HR. It differs from the PÉGASE.2’s color code only by the input format of the data (FITS format instead of ASCII format). Run for example:

    > colors_HR myspectrum.fits

to generate an ASCII file containing the evolution of many colors.

colors HR makes use of the files data/user defined/filters.dat and data/user defined/calib.dat.

This calib.dat file is created automatically by the initial make command, with the code bin/calib HR and from the filters.dat file (see PÉGASE.2’s README for the details of this calibration).

5 References

Fioc M., Rocca-Volmerange B., 1997, A&A 326, 950 Fioc M., Rocca-Volmerange B., 2000, astro-ph/9912179 (PÉGASE.2’s README) Katz D., Soubiran C., Cayrel R., Adda, M., Cautain R., 1998, A&A,338,151 Le Borgne D., Rocca-Volmerange B., Prugniel Ph., Lançon A., Fioc M., Soubiran C., 2004, A&A in

press, astro-ph/0408xxx Prugniel Ph., Soubiran C., 2001, A&A,369,1048 Soubiran C., Katz D., Cayrel R., 1998, A&AS,133,221